OpenStudy (anonymous):

write an equation point-slope form for the line through the given point with the given slope. (-4,-2); m= - 1/2

OpenStudy (anonymous):

do you remember what point slope form is? \[y-y=m(x-x)\]

OpenStudy (anonymous):

no im not sure i totally forgot @lonnie455rich

OpenStudy (anonymous):

well. the form I posted is point slope. you know that m = slope right?

OpenStudy (anonymous):

nooooo

OpenStudy (anonymous):

well it is. knowing all this, take a stab at it and ill let you know if you have any mistakes

OpenStudy (anonymous):

y+7=-0.54(x-4)

OpenStudy (anonymous):

@lonnie455rich

OpenStudy (anonymous):

where did you get those numbers from. I see (-4,-2) as points and slope is -1/2

OpenStudy (anonymous):

i was doing the wrong question okay i got y+2=-1/2(x-4)

OpenStudy (anonymous):

is it right

OpenStudy (anonymous):

almost. think about your sign on the (x-4) you got it correct on the left side with the y. why didn't you on the right?

OpenStudy (anonymous):

+ sign ?

OpenStudy (anonymous):

yes because if you write it all out you get \[y-(-2)=-1/2(x-(-4)\] then if you simplify what would you get

OpenStudy (anonymous):

y+2=-1/2(x+4)

OpenStudy (anonymous):

perfect.

OpenStudy (anonymous):

thanksss can you help with some more ?

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

write an equation point-slope form for the line through the given point with the given slope. (4-7);m=-0.54 my answer = y-7==0.54(x-4) right ?

OpenStudy (anonymous):

with the points (4,-7) and m=-.54 looks like I see 1-2 mistakes depending on if one of those equal signs is supposed to be a negative sign. think about the -7 and what we did earlier

OpenStudy (anonymous):

okay so it would be y+7=-0.54(x+4)

OpenStudy (anonymous):

well you made a mistake there. think about your sign on 4

OpenStudy (anonymous):

- right ? i always get my signs confused

OpenStudy (anonymous):

it will only be positive there if your point is negative because of the point slope formula

OpenStudy (anonymous):

whenever I learn any kind of new technique or equation. on every problem I do. I just write Eq: then write out the equation then I solve the problem. helps me remember things.

OpenStudy (anonymous):

okay thanks thats better because i get confused a lot and im not so good in math you know....

OpenStudy (anonymous):

math is arguably 1st-2nd of the hardest subjects there is for any given level of schooling. don't worry, just takes practice.

OpenStudy (anonymous):

thank you for helping me i understand it better now ( : yes im going to pratice now

OpenStudy (anonymous):

glad to hear. good luck